OpenStudy (anonymous):

An arithmetic sequence is defined by the general term tn = -5 + (n - 1)78, where n ∈N and n ≥ 1. What is the recursive formula of the sequence?

OpenStudy (ranga):

\[\Large t_n = -5 + (n-1)78\]\[\Large t_{n+1} = -5 + ((n+1)-1)78\]\[\Large t_{n+1} - t_n = -5 + ((n+1)-1 - n + 1)78 = 78\]\[\Large t_{n+1} = t_n + 78\]

OpenStudy (ranga):

third line should have - 5 + 5.

OpenStudy (anonymous):

thank you so much! this helped a lot

OpenStudy (ranga):

you are welcome.